OpenStudy (anonymous):

The length of a rectangular room is 7.9 m and its width is 8.6 m. Find the area of the room A. 73.96 m2 B. 62.41 m2 C. 67.94 m2 D. 33 m2

OpenStudy (anonymous):

Just multiply: \[7.9\times8.6\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have another qustion

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You cut square corners with side lengths that are whole numbers from a piece of cardboard with dimensions 20 inches by 30 inches. You then fold the cardboard to create a box with no lid. Which of the following dimensions will give you the greatest volume?

OpenStudy (anonymous):

A. 12 in. by 22 in. by 4 in. B. 10 in. by 20 in. by 5 in. C. 14 in. by 24 in. by 2 in. D. 10 in. by 24 in. by 6 in

OpenStudy (anonymous):

A particular model of walkie-talkie can broadcast in a circular area. The radius of the broadcast area is 10,000 feet. Find the area of this circle to the nearest square foot. Use 3.14 for p. A. 314,000,000 ft2 B. 100,000,000 ft2 C. 1,256,000,000 ft2 D. 62,800 ft2

OpenStudy (anonymous):

On the first, ignore the word problem, it's just trying to confuse you. Take each of the answers, and multiply them together (e.g. 12 x 22 x 4) to get the volume of each and see which is biggest.

OpenStudy (anonymous):

D. 10 in. by 24 in. by 6 in ?

OpenStudy (anonymous):

The area of a circle is \[pi \times r ^{2}\] So, you would need to calculate \[3.14 \times 10,000 \times 10,000\]

OpenStudy (anonymous):

ok i got it :) Find the volume of a can of soup that has a height of 16 cm and a radius of 5 cm. Use 3.14 for p. A. 1,256.0 cm3 B. 251.2 cm3 C. 4,019.2 cm3 D. 502.4 cm3

OpenStudy (anonymous):

Yes to D above. Again, find the area of the circle \[pi \times5\] And then multiply it by the height of the can (16) to get the volume of the cylinder. \[pi \times 5 \times 16\]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Find the volume of the sphere to the nearest whole number. Use p = 3.14. the radius is 5in A. 131 in.3 B. 393 in.3 C. 4,187 in.3 D. 523 in.3

OpenStudy (anonymous):

The volume of a sphere is \[(4 \div 3) pi \times r ^{3}\] So \[(4/3) \times 3.14 \times 5 \times 5 \times 5\]

OpenStudy (anonymous):

i need u to be my tutor, ur sooo smart

OpenStudy (anonymous):

You are designing a new container for powdered laundry detergent. You are considering a cylindrical container with a diameter of 14 inches and a height of 18 inches. Find the volume of this container to the nearest cubic unit. Use a calculator. A. 1,100 in.3 B. 882 in.3 C. 11,084 in.3 D. 2,771 in.3

OpenStudy (anonymous):

Write the most precise name for the space figure with the given properties. a lateral surface and two circular bases A. prism B. sphere C. cone D. cylinder

OpenStudy (anonymous):

This is a cylinder again, so use \[pi \times r ^{2} \times height\] So \[pi \times 14 ^{2} \times 18\] It tells you to use a calculator, because it wants a precise value for pi (use 3.1416 instead of 3.14).

OpenStudy (anonymous):

A. 124 cm2 B. 110 cm2 C. 150 cm2 D. 164 cm2

OpenStudy (anonymous):

i used the formula but i came up with 11,077

OpenStudy (anonymous):

oh i got 11,083 which is C 11,084 in.3